OpenStudy (anonymous):

can anyone help me? @sheelo_mughal @texaschic101 @YanaSidlinskiy @mathmale @ankitshaw

OpenStudy (anonymous):

I can!

OpenStudy (anonymous):

for what just type ur ques.

OpenStudy (anonymous):

Solve the inequality 2x + 8 < 5x - 4. I will show you what i got so far

OpenStudy (anonymous):

This is what i got, anyone else agree? 2x+8<5x-4 10x2<x-4 5x<4 x>4

OpenStudy (anonymous):

I got x<4 what did i do wrong?

OpenStudy (anonymous):

You didn't reverse the sign at the end of the problem

OpenStudy (anonymous):

You kept it like this <, But you were suppose to change it to this >

OpenStudy (anonymous):

why?

OpenStudy (texaschic101):

2x + 8 < 5x - 4 2x - 5x < -4 - 8 -3x < - 12 -- when dividing/multiplying by a negative, the sign changes x > 4

OpenStudy (anonymous):

@texaschic101 Bingo!

OpenStudy (anonymous):

4x-2x-2=3x+10 2x-2=3x+10 2x-3x=10+2 -x = 12 x= -12 This was next problem

OpenStudy (texaschic101):

correct

OpenStudy (anonymous):

Your correct!

OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 6 - x = 5x + 30 4x - 8 = 3x + 2 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

this one i dont understand at all

OpenStudy (anonymous):

i was thinking A but im not 100% sure

OpenStudy (texaschic101):

6 - x = 5x + 30 -x - 5x = 30 - 6 -6x = 24 x = -4 4x - 8 = 3x + 2 4x - 3x = 2 + 8 x = 10 solution to equation 1 is smaller then equation 2

OpenStudy (anonymous):

Man! You beat me to it! @texaschic101

OpenStudy (texaschic101):

lol....sorry. I didn't mean to take over your question

OpenStudy (anonymous):

I know it seems like i dont need help because so far everything has been pretty much correct but i like to make sure the work i do is 100% correct. So thank you both.

OpenStudy (anonymous):

Lol, its cool

OpenStudy (anonymous):

Anytime! @jadee27p

OpenStudy (texaschic101):

jadee...confirmation is always good.

OpenStudy (anonymous):

I have 2 more. Solve for x: -3|x - 3| = -6 x = 1, x = -1 x = 1, x = 5 x = 0, x = 5 No solutions

OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 -5|3x - 2| = -15 |2x - 9| = 11 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to equations shown.

OpenStudy (anonymous):

and that's all.

OpenStudy (anonymous):

-5|3x - 2| = -15 has two solutions= -1/3 and 5/3 |2x - 9| = 11= Also has two solutions= -1 and 10

OpenStudy (anonymous):

Therefore i would go with, [Equation #1 and Equation #2 have the same number of solutions.]

OpenStudy (anonymous):

Solve for x: -3|x - 3| = -6 x = 1, x = -1 x = 1 x = 5 x = 0 x = 5 No solutions What about this one

OpenStudy (anonymous):

Are they your answer choices or the answer you got?

OpenStudy (anonymous):

Choices

OpenStudy (anonymous):

Okay :)

OpenStudy (texaschic101):

-3 | x - 3| = -6 --- divide both sides by -3 | x - 3 | = 2 x - 3 = 2 -(x - 3) = 2 x = 2 + 3 -x + 3 = 2 x = 5 -x = 2 - 3 -x = -1 x = 1 xolutions are : x = 5 and x = 1 um...I am getting this

OpenStudy (anonymous):

For that one i would go with x=1 and/or x=5

OpenStudy (anonymous):

In order for solving problems like this what would i start off with ?

OpenStudy (texaschic101):

get just what is in the absolute value on one side, and everything else on the other side

OpenStudy (anonymous):

Thank you guys :)

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

No problem!